# references.bib

### v1

@@ Abstract
Regulatory functions are essential in both socioeconomic and biological systems, from corporate managers to regulatory genes in genomes. Regulatory functions come with substantial costs, but are often taken for granted. Here, we empirically examine regulatory costs across diverse systems—biological organisms (bacteria and eukaryotic genomes), human organizations (companies, federal agencies, universities), and decentralized entities (Wikipedia, cities)—using scaling analysis. We guide the empirical analysis with a conceptual model, which anticipates the scaling of regulatory costs to shift with the system’s internal interaction structure—well-mixed or modular. We find diverse systems exhibit consistent scaling patterns—well-mixed systems exhibit superlinear scaling, while modular ones show sublinear or linear scaling. Further, we find that the socioeconomic systems containing more diverse occupational functions tend to have more regulatory costs than expected from their size, confirming the type of interactions also plays a role in regulatory costs. While many socioeconomic systems exhibit efficiencies of scale, regulatory costs in many social systems have grown disproportionally over time. Our finding suggests that the increasing complexity of functions may contribute to this trend. This cross-system comparison offers a framework for understanding regulatory costs and could guide future efforts to identify and mitigate regulatory inefficiencies.
@@ 1 Introduction
Regulatory functions and mechanisms are necessary and ubiquitous features across all biological and social complex adaptive systems, spanning scales from the cellular to societal level. Regardless of scale, organizational structure, or composition, these systems rely on dedicated components to regulate internal processes and mediate interactions among their constituent parts. For instance, in cells, some genes may be beneficial when expressed in isolation, but their expressed proteins can often interact adversely. Regulatory genes control the timing of gene expression, preventing these harmful interactions by ensuring that conflicting genes are not expressed simultaneously. Similarly, in companies, managers coordinate employee activities to prevent duplicated efforts and mitigate adverse interpersonal interactions. In broader societal contexts, legal professionals mediate interactions among citizens through legal processes, ensuring orderly conduct.
@@
While necessary and ubiquitous, regulatory functions consume a significant amount of energy and resources. For example, regulatory genes in bacteria account for about $10$% of the costs. In the US, 15% of workforce compensation is paid to managers [blsWages]. In US universities, administrative spending is on par with instructional spending, reaching \$122.3 billion in the 2014–15 school year, and has been cited as a key factor in the skyrocketing tuition of US universities [mumper2005causes, greene2010administrative]. In fact, the burden of administrative costs is a significant concern in many aspects of society, including higher education [ginsberg2011fall], health care [himmelstein2020health], manufacturing [meyer2013limits], and the transition to renewable technologies [klemun2023mechanisms], As a result, regulatory costs have emerged as one of the major societal challenges of the 21st century. What aspects of regulatory function are set by fundamental requirements and which are malleable through changes in structure, culture, or procedure? The answers to these larger questions require more understanding of the mechanisms underlying regulatory functions across a wide range of systems.
Regulatory costs have been studied separately at the levels of organizations, societies, and biological organisms. In socioeconomic systems, the primary focus has been on best practices for management and regulations concerning profits and efficiencies [VanReenen2010JEP], while in biology, the emphasis has been on understanding the genetic and metabolic mechanisms of regulation such as complex feedback loops and signaling pathways that maintain cellular homeostasis and respond to environmental changes. These functions and expenses seem to correspond to the expenses related to governance, law enforcement, and social coordination to regulate and coordinate activities, monitor performance, and enforce rules within the organization [yoon2023makes]. As such, studies have identified several key determinants of regulatory costs, including size, measured by the number of employees [boyne2013burdened], internal structure, measured by the level of hierarchies and size of the sub-units [carillo1991organization], and functional complexity [anderson1961organizational], often measured by the number of different tasks performed by individuals within the organization. Despite these valuable insights, a comprehensive and systematic understanding of the determinants of regulatory costs remains elusive.
@@
In the study of biological organisms, we have a much clearer understanding of the fundamental and baseline requirements for regulatory function. A key assumption in the biological context is that the overall evolutionary process, involving vast species and timescales, leads to optimized regulatory costs. Thus, biological examples provide a case study for understanding essential or optimal regulatory functions. Furthermore, similar key determinants of size, internal structure, and complexity emerge from the biological literature. For instance, major transitions in biological architecture—–such as the evolution from bacteria, which lack internal compartments and allow any gene to interact with any other gene, to eukaryotes, which have compartmentalized cells, and from single-cell to multicellular organisms—each introduce new and distinct forms of regulation [szathmary2015toward]. At a more detailed level, all cellular functions, including regulatory genes, are known to scale with organism size [kempes2016evolutionary], [kempes2012growth, gondhalekar2023scaling, van2003scaling].
@@
The identification of common key determinants—size, structure, and functional complexity—across both social and biological systems suggests the potential for a unifying and comparative perspective. Previous studies have successfully achieved this goal by using scaling analysis as a foundation for empirical analysis. By combining this with mechanistic models that explain scaling relationships, researchers have uncovered common mechanisms across a diverse range of systems, including those in physics, biology, ecology, as well as firms and cities [west2018scale, van2003scaling, kempes2017drivers, cody1975ecology, bettencourt2007growth]. These studies reveal key connections between size, function, and architecture, illustrating how fundamental principles can be applied across different types of systems.
@@
Here, we first conceptualize what gives rise to regulatory costs across complex systems based on managing adverse interactions by integrating the interaction of size, structure, and various kinds of costs. We compare and unify diverse systems across the dimensions of well-mixed to highly modular. At one end of the spectrum, well-mixed systems tend to be the simplest self-organized and agglomerate ones. At the other end, those with modular structures tend to be centrally planned or have gone through several transitions in architecture. For example, bacteria are defined by a cellular environment where most expressed proteins in the liquid cytoplasm can diffuse and interact with any other expressed protein, posing unique regulatory challenges for governing protein co-expression. The growth in genome size for bacteria is a result of an emergent process of evolution. In contrast, companies are typically defined by hierarchical and modular structures that regulate interactions amongst individuals and are highly planned as companies grow in size. This pattern also applies to more complex organisms that have experienced multiple major evolutionary transitions in structure [szathmary2015toward]. We then compile data on regulatory costs in biological and social systems, including regulatory genes in bacteria and eukaryote cells, managers in companies, governmental agencies, universities, administrators on Wikipedia, and lawyers in cities. Using scaling analysis, we quantify shifts in regulatory costs with entity size by the scaling exponents to reveal shared governing processes and principles across different systems.
@@ 2 Results
We define regulatory functions as entities whose primary role is to moderate, adjust, or coordinate the interactions among other entities. These include regulatory genes in cells, managers in companies, legal functions in cities, and administrators on Wikipedia. Examples of functions that do not fall in this category include primarily functional components, for example, functional genes in cells, factory workers in a manufacturing plant, and primarily maintenance and repair functions, such as janitors of a university.
2.1 Conceptualizing the mechanisms for baseline regulatory costs across systems
@@
In this section, we present a simple mathematical framework to illustrate how baseline requirements for regulation may be derived from the interactions of a few key mechanisms. The central premise is that regulation arises as a response to potential adverse interactions among system components [malone1994coordination]. A complex system brings together a large number of individual components—such as genes in cells or employees in companies—to achieve certain benefits, like metabolic energy for cells or revenue for companies. We denote the amount of these benefits as $B$. However, adverse interactions among these components can diminish these benefits. In cells, such interactions may occur when expressed proteins interact in ways that are futile or detrimental to cellular metabolism. In organizations, adverse interactions can manifest as duplicated efforts or interpersonal conflicts between individuals. When two components of a system have an adverse interaction, we identify three responses the system can employ, as illustrated in Fig. 1. First, the system can do nothing and tolerate the adverse interaction. Second, it can use a regulator to manage these interactions. In cells, it takes the form of carrying a regulatory gene, which makes sure the genes encoding two negatively interacting proteins are not expressed at the same time. In organizations, it can take the form of assigning a manager to coordinate tasks and prevent duplicated work between individuals or mediate interpersonal conflict between two employees. Third, the system can separate the two components by creating compartments. In cells, this involves developing internal architecture, such as mitochondria, to ensure that certain genes are only expressed in specific sub-portions of the cell. In organizations, this approach could involve structuring individuals into separate teams, units or departments, thereby modularizing their efforts.
@@
Each of these three strategies carries a cost. esebreak three categories as: the cost of adverse interactions, $I$, the cost associated with regulators in the system, $R$, and the cost associated with compartments $C$. Regulators and compartments each reduce adverse interactions, but come with their own costs such as the pay required to employ a manager or the energy dedicated to maintaining and expressing a regulatory gene. Compartments differ from regulators in that they use structural separation, such as organizational divisions or sub-cellular membranes, to isolate unnecessary interactions. It’s important to note that the cost associated with regulators and compartments may interact. For instance, in companies, the establishment of a new compartment, like a new division, is typically accompanied by the appointment of a regulator, such as the division head.
@@
We are interested in how these benefits and costs change with the system size, $N$, such as the number of employees in an organization or the number of genes in a genome. In particular, we specify the number of functional individuals, $f$, regulators, $r$, and the number of compartments, $c$, to arrive at the generic utility function
@@
Equation 1 enables the optimization of the utility function, given functional forms for the four terms over the number of regulators, $r$, and number of compartments, $c$. While the precise quantitative forms of these four terms still remain uncertain for specific systems, it is useful to define the simplest version of each term to understand the baseline optimization of $\mathcal{L}$ and how $r$ and $c$ scale with system size.
In a well-mixed environment, the addition of a new functional individual—whether an employee or gene—comes with a probability $\mu$ of having a negative interaction with all existing functional members of the system. Denoting the number of functional individuals as $f$, edge counting gives the total number of negative interactions as $\rho f\left(f-1\right)$, with $\rho\equiv\mu/2$. Each of these negative interactions comes with a cost $\gamma_{1}$, and we can remove negative interactions either by adding regulators or placing individuals in compartments. We assume that individuals contained within compartments are removed from pairwise negative interactions with the rest of the system. If we place $\eta$ individuals into a compartment, and the system contains $c$ compartments, then the total number of negative interactions becomes $\rho\left(f-\eta c\right)\left(f-\eta c-1\right)$. For a fixed compartment size, negative interactions within a compartment simply become a fixed cost, which we will combine with compartment costs later. Additionally, if each regulator can reduce $\theta$ negative interactions, then the total number of negative interactions becomes $\rho\left(f-\eta c\right)\left(f-\eta c-1\right)-\theta r$ leading to the final utility function,
@@
where $\gamma_{2}$ is the unit cost of a regulator and $\gamma_{3}$ is the unit cost of a compartment (including the cost of negative interactions and regulators within a compartment). Here we assume that there is a proportional cost of $r$ and $c$ and a linear increase in benefit associated with $f$ such that $R=\gamma_{2}r$, $C=\gamma_{3}c$, and $B=bf$. All terms are measured in dollars or energy depending on the system of interest. It is also useful to note that the total size of the system, $N$, which is often what is measured, is the sum of functional and regulatory components, $N=f+r$.
@@
We now optimize $\mathcal{L}$ as a function of $r$ and $c$. Later we will show that compartments are not beneficial below a certain size, and so we begin by considering $\mathcal{L}$ with $c=0$. From Equation 2, we see that regulators are only beneficial if $\gamma_{1}>\gamma_{2}/\theta$. That is, the cost of regulating a negative interaction is less than the cost of the negative interaction itself.
The utility function is optimized when regulators mediate all adverse interactions, meaning $I\left(f,r\right)=0$. This gives the optimal number of regulators,
$r_{opt}=\frac{f(f-1)\rho}{\theta}\sim\frac{f^{2}\rho}{\theta}$ (3)
@@
illustrating that $r$ grows like $f^{2}$ adjusted by the ratio of the probability of negative interactions, $\rho$ to the number of interactions that one regulator can mitigate $\theta$. Given that $r\sim f^{2}$ we can also write the total size in terms of $r$ alone where $N=f+r=\sqrt{\theta/\rho}\;r^{1/2}+r$. It is useful to note that we have two regimes here: $N\propto r^{1/2}$ for small $f$ (which is also small $r$ given the relationships above) and $N\propto r$ in the large $f$ limit.
The first case compares well to bacteria where every expressed gene can interact with every other gene. Here the observed scaling is $r\sim N^{1.86}$ [van2003scaling], in rough agreement with this simple optimization where $r\propto N^{2}$.
For the second case, where $N\propto r$, there is a critical size at which regulators overwhelm the system causing $\mathcal{L}=0$. For the optimal number of regulators, where $I=0$, the utility function is
$\mathcal{L}=bf-\frac{\gamma_{2}\rho f\left(f-1\right)}{\theta}.$ (4)
which equals zero at
$f_{r}=1+\frac{b\theta}{\gamma_{2}\rho}.$ (5)
@@
Here $f_{r}$ is the maximum number of functional elements for a system with only regulators. This upper bound depends on the ratio of the benefit per functional element to an effective cost, the unit cost of a regulator times the likelihood of a negative interaction. The ratio of $b\theta$ to $\gamma_{2}\rho$ could be quite large given that the probability of negative interactions could be small and the cost of an average individual, including regulators, $\gamma_{2}$, should be small relative to the productive output, $b$, of an average individual. In addition, a regulator may be able to handle many interactions and $\theta$ should be much larger than 1. For example, if $\rho=0.10$, $\gamma_{2}=b/2$, and $\theta=10$, then $f\approx 200$, implying that an organization that handles negative interactions with only regulators could reach significant size.
@@
As we mentioned earlier, $f_{r}$ defines the size at which regulators become the entire system and this implies the need for compartments. Optimizing the complete cost function of Equation 2 with respect to $c$ by setting $\partial\mathcal{L}/\partial c=0$ yields the optimal number of compartments,
$c_{opt}=\frac{1}{2\eta}\left(2f-1-\frac{\gamma_{3}}{\gamma_{1}\rho\eta}\right)$ (6)
This result shows that compartments are only advantageous for cells with a large enough number of functional genes since $c_{opt}>0$ requires
$f_{c}>\frac{1}{2}\left(1+\frac{\gamma_{3}}{\gamma_{1}\rho\eta}\right).$ (7)
The critical value is set by the ratio of the unit cost of a compartment to the probability of negative interactions multiplied by the unit cost of a negative interaction and how many elements are put in a compartment. In this case, the unit cost of a compartment, denoted as $\gamma_{3}$, should exceed the unit cost of negative interaction. This is because each compartment may entail the cost of physical space, regulators who maintain it, and coordination costs among compartments. The scale at which regulators become the entire system, $f_{r}$, will agree with the minimum size that compartments are feasible, $f_{c}$. The criterion for a smooth transition to compartments is thus given by $f_{r}=f_{c}$, which sets the following condition for the unit costs
$\gamma_{3}=\gamma_{1}\eta\left(\rho+\frac{2b\theta}{\gamma_{2}}\right).$ (8)
To give a sense of the size of the unit compartment costs, $\gamma_{3}=40\gamma_{1}\eta$ leads to the same transition point of $f_{c}\approx 200$ for the values of $b$ and $\rho$ used above.
Above the critical value $f_{c}$, the number of compartments scales like $c\sim f$, which also affects the scaling of the number of regulators. Given the optimal number of compartments, we can maximize the utility function by considering the point at which $\mathcal{L}=bf$, which occurs at
$r_{opt}=\frac{c\gamma_{3}+\gamma_{1}\rho(f-\eta c)(f-\eta c-1)}{\gamma_{1}\theta-\gamma_{2}}$ (9)
which given the solution for $c_{opt}$ becomes
$r_{opt}=\frac{4\gamma_{3}f-\gamma_{1}\eta\rho-2\gamma_{3}-\gamma_{3}^{2}/\left(\gamma_{1}\eta\rho\right)}{4\eta\left(\gamma_{1}\theta-\gamma_{2}\right)}$ (10)
which for large $f$ is approximated by
$r_{opt}\approx\frac{\gamma_{3}}{\eta\left(\gamma_{1}\theta\ -\gamma_{2}\right)}f.$ (11)
@@
With compartments, these results show that $r\sim f$ and $c\sim f$. Because of this linear scaling, the total $N=f+r$ is also linear $r\propto N$ and $c\propto N$. In this regime, quadratic requirements of negative pairwise interactions are proportionally handled by compartments and regulatory genes. However, this only occurs after the transition where compartments become inexpensive enough to be viable.

### v2

@@ Abstract
Regulatory functions are essential in both socioeconomic and biological systems, from corporate managers to regulatory genes. Regulatory functions come with substantial costs and benefits, and the balance of the two is often taken for granted. A fundamental question for all complex systems becomes how much regulatory function do they need for their size and function? Here, we present empirical evidence that regulatory functions scale systematically across diverse systems: biological organisms (bacterial and eukaryotic genomes), human organizations (companies, federal agencies, universities), and decentralized entities (Wikipedia, cities). We combine an analysis of large data sets from each of these domains with a simple conceptual model. The model predicts that the scaling of regulatory costs shifts with system structure. Well-mixed small systems exhibit superlinear scaling between size and regulatory function, while modular large ones show sublinear or linear scaling, both in agreement with data. Finally, we find that socioeconomic systems that contain more diverse occupational functions tend to have more regulatory costs than expected from the scaling relationships, confirming the hypothesis that the type and complexity of interactions also play a role in regulatory costs. Our cross-system comparison offers a mechanistic framework for understanding regulatory function and can potentially guide efforts to analyze the costs and benefits of regulatory function in diverse systems.
@@ 1 Introduction
Regulatory functions and mechanisms are fundamental for maintaining stability, managing complexity, and coordinating interactions in both biological and social complex adaptive systems. These processes operate across different scales, from individual cells to entire societies, and rely on dedicated components that regulate internal processes and mediate interactions among constituent parts. In biological systems, regulatory functions prevent harmful interactions between components that may be individually beneficial but detrimental when expressed together. For this, regulatory genes control the timing and conditions of gene expression, ensuring that conflicting interactions are avoided. For example, allosteric regulation of proteins, such as serpins, helps inhibit diseases such as emphysema [sanrattana2019serpins]. Likewise, in human organizations, regulation is presumed to optimize efficiency and prevent dysfunction by controlling workflows, overseeing tasks, and resolving conflicts [mitzberg1973, mitzberg1979]. At the societal level, legal and institutional regulations mediate disputes, allocate resources, and prevent systemic failures and thus maintain order and balance among competing interests [weber1978max].
@@
These necessary and ubiquitous regulatory functions consume a significant amount of energy and resources. For example, regulatory genes in bacteria account for about $10$% of the metabolic costs [lynch2015bioenergetic]. In the US, 15% of workforce compensation is paid to managers [blsWages]. In US universities, administrative spending is on par with instructional spending, and has been cited as a key factor in the skyrocketing tuition cost [mumper2005causes, greene2010administrative]. In fact, the burden of administrative costs is a significant concern in many aspects of society, including higher education [ginsberg2011fall], health care [himmelstein2020health], manufacturing [meyer2013limits], and the transition to renewable technologies [klemun2023mechanisms]. As a result, regulatory costs have emerged as one of the major societal challenges of the 21st century. What aspects of regulatory function are set by fundamental requirements and which are malleable through changes in structure, culture, or procedure? The answers to these big questions require more understanding of the mechanisms underlying regulatory functions across a wide range of systems, but a unified understanding across these domains has yet to be developed.
@@
Historically, regulatory costs have been studied independently in the domains of biological organisms, organizations, and societies, yet a comprehensive understanding that bridges these domains has remained a longstanding open challenge. In the study of biological organisms, we have a clearer understanding of the fundamental and baseline requirements for regulatory function. A key assumption in the biological context is that the overall evolutionary process, involving vast numbers of species and multiple timescales, tends towards the optimization of regulatory costs [kempes2019scales]. Thus, biological examples provide a useful case study for understanding essential or optimal regulatory functions. Furthermore, similar key determinants of size, internal structure, and complexity emerge from the biological literature. For instance, major transitions in biological architecture — such as the evolution from bacteria, which lack internal compartments and allow any gene to interact with any other gene, to eukaryotes, which have compartmentalized cells and multiple chromosomes, to multicellular organizations — each introduces new and distinct forms of regulation [szathmary2015toward].
Similarly, regulatory costs in socioeconomic systems are highly complex and take many different forms, each representing an area of study, including managerial overhead, administrative intensity, and bureaucratic burdens. For example, in corporate environments, the role of managers has been extensively studied, with particular attention to best practices for improving efficiency and coordination [VanReenen2010JEP]. Middle and upper management oversee workflows, resolve conflicts, and align operations with institutional goals, yet these functions also contribute to administrative overhead [VanReenen2010JEP, Hjort2022, Chandler1977]. Similarly, regulatory structures in government and other institutions mediate interactions, allocate resources, and prevent systemic failures, ensuring stability within complex social systems [jennings2005weber, hannan1984structural, Ostrom].
Across these diverse contexts, regulatory mechanisms in both biological and social systems share a fundamental purpose: maintaining order and preventing or resolving conflicts [malone1994coordination, herbert1947]. In biological systems, regulation operates through genetic and metabolic mechanisms that maintain cellular homeostasis and enable adaptive responses to environmental changes [konieczny2023regulation]. In human organizations, regulatory functions manifest through governance, law enforcement, and institutional coordination, ensuring social stability and managing disputes [march1993organizations, mintzberg1980structure, herbert1947, Ostrom]. In both cases, regulation serves to monitor activities, enforce rules, and maintain overall system stability.
There are several key determinants of regulatory costs across these domains. These include size, often measured by the number of individuals in the system, such as employees or genomes [boyne2013burdened, van2003scaling]; internal structure, measured by the level of hierarchies and size of the sub-units [carillo1991organization]; and functional complexity [anderson1961organizational], often measured by the number of different tasks performed by individuals. While these factors are well-documented within their respective fields, a comprehensive and systematic understanding of the determinants of regulatory costs remains elusive.
@@
The identification of common key determinants across both social and biological systems, such as size, structure, and functional complexity, suggests the potential for a unifying and comparative perspective. A powerful tool that has been successful in accomplishing this across a diverse range of systems, including those in physics, biology, and ecology, as well as for firms and cities [west2018scale, van2003scaling, kempes2017drivers, cody1975ecology, bettencourt2007growth] is the use of scaling analyses to reveal common underlying mechanisms. Fundamental metrics in many of these systems exhibit simple power law scaling behavior as a function of their size, which can be mechanistically, and in many cases quantitatively, understood from similar underlying dynamics. Such studies reveal key connections between size, function, and architecture, illustrating how fundamental principles can be applied across different types of systems. A major contribution of this paper is to show that metrics reflecting regulatory functions from bacteria to cities and companies do indeed scale in a similar and systematic fashion, strongly suggesting that there are unifying principles and dynamics at play.
@@
First, we conceptualize what gives rise to regulatory costs across complex systems based on the costs to manage adverse interactions. We compare, contrast and unify the degree to which the constituents of these diverse systems are well-mixed or highly modular. At one end of the spectrum, well-mixed systems tend to be the simplest self-organized and agglomerated ones. At the other end, those with modular structures tend to be centrally planned or have gone through several transitions in architecture. For example, bacteria are defined by a cellular environment where most expressed proteins in the liquid cytoplasm can diffuse and interact with any other expressed protein, posing unique regulatory challenges for governing protein co-expression. At the genome level, in bacteria the genome is physically distributed in a nucleoid with no surrounding compartment, in contrast to the genome of eukaryotes which is located in the nucleus with a complex compartmentalization, that includes not just distinct chromosomes, but also territories and other structures [meldi2011compartmentalization, cremer2010chromosome]. In contrast to bacteria, companies are typically defined by hierarchical and modular structures that regulate interactions among individuals and are highly planned as they grow in size. This pattern also applies to more complex organisms that have experienced multiple major evolutionary transitions in structure [szathmary2015toward]. We then compile data on regulatory costs in biological and social systems, including regulatory genes in both bacteria and eukaryote cells, managers in companies, governmental agencies and universities, administrators on Wikipedia, and lawyers in cities. We show that all of these scale with system size as power laws, indicating that, despite the broad diversity, they are all manifestations of self-similar structures. Their apparent differences can be quantified by the value of the scaling exponents in how their regulatory costs change with size to reveal shared governing processes and principles across different systems.
@@ 2 Results
We define regulatory functions as entities whose primary role is to moderate, adjust, or coordinate the interactions among other entities. These include regulatory genes in cells, managers in companies, lawyers in cities, and administrators on Wikipedia. Examples of functions that do not fall in this category include primarily functional components, for example, functional genes in cells, factory workers in a manufacturing plant, and primarily maintenance and repair functions, such as janitors of a university.
2.1 Model for baseline expectations of regulatory costs across systems
@@
We present a simple mathematical framework to illustrate how baseline requirements for regulation can be derived from the interactions of a few key mechanisms. The central premise is that regulation arises as a response to potential adverse interactions among system components [malone1994coordination]. A complex system brings together a large number of individual components — such as genes in cells or employees in companies — to achieve certain benefits, like metabolic energy for cells or revenue for companies. We denote the amount of these benefits as $B$. However, adverse interactions among these components can diminish these benefits. In cells, such interactions may occur when expressed proteins interact in ways that are futile or detrimental to cellular metabolism. In organizations, adverse interactions can manifest as duplicated efforts or interpersonal conflicts between individuals. When two components of a system have an adverse interaction, we identify three responses the system can employ, as illustrated in Fig. 1. First, the system can do nothing and tolerate the adverse interaction. Second, it can use a regulator to manage these interactions. In cells, it takes the form of carrying a regulatory gene, which makes sure the genes encoding two negatively interacting proteins are not expressed at the same time. In organizations, it can take the form of assigning a manager to coordinate tasks and prevent duplicated work between individuals or mediate interpersonal conflict between two employees. Third, the system can separate the two components by creating compartments. In cells, this involves developing internal architecture, such as mitochondria, to ensure that certain genes are only expressed in specific sub-portions of the cell. In organizations, this approach could involve structuring individuals into separate teams, units or departments, thereby modularizing their efforts.
@@
Each of these three strategies carries a cost. We denote costs in these three categories as: the cost of adverse interactions, $I$, the cost associated with regulators in the system, $R$, and the cost associated with compartments $C$. Regulators and compartments each reduce adverse interactions, but come with their own costs such as the pay required to employ a manager or the energy dedicated to maintaining and expressing a regulatory gene. Compartments differ from regulators in that they use structural separation, such as organizational divisions, sub-cellular membranes, or genomic regions, to isolate unnecessary interactions. It’s important to note that the cost associated with regulators and compartments may interact. For instance, in companies, the establishment of a new compartment, like a new division, is typically accompanied by the appointment of a regulator, such as the division head.
Figure 2: The scale-based model of function, regulation, and compartments. Functional elements are able to freely interact without compartments or regulators. Regulators modulate pairwise interactions and compartments restrict interactions to functional elements within the same compartment. The size of the system, $N$, is the total number of functional elements. Equation 1 quantifies the costs associated with negative interactions, regulators, and compartments.
@@
We are interested in how these benefits and costs change with the system size, $N$, such as the number of employees in an organization or the number of genes in a genome (Fig. 2). In particular, we specify the number of functional individuals, $f$, regulators, $r$, and the number of compartments, $c$, to arrive at the generic utility function
@@
Equation  1 enables the optimization of the utility function, given the analytical forms for the four terms, $B,I,R$ and $C$, as functions of the number of regulators, $r$, and the number of compartments, $c$. While the precise quantitative forms of these four terms still remain uncertain for specific systems, it is useful to introduce the simplest version of each term to understand the baseline optimization of $\mathcal{L}$ and how $r$ and $c$ scale with system size. We present a summary of our findings below, while the full derivation can be found in the Supplementary Information.
In a well-mixed environment in which all constituent components interact equally with all others, the addition of a new functional individual—whether an employee or gene—comes with a probability $\mu$ of having a negative interaction with all existing functional members of the system. If we consider the $f$ functional individuals as identical, then the total number of negative pairwise interactions is $\mu f\left(f-1\right)/2$. Each of these negative interactions comes with a cost $\gamma_{1}$, and we can remove negative interactions either by adding regulators or placing individuals in compartments. We assume that individuals contained within compartments are removed from pairwise negative interactions with the rest of the system. If we place $\eta$ individuals into a compartment, and the system contains $c$ identical compartments, then the total number of negative interactions becomes $\rho\left(f-\eta c\right)\left(f-\eta c-1\right)$, where $\rho\equiv\mu/2$. For a fixed compartment size, negative interactions within a compartment simply become a fixed cost, which we will combine with compartment costs later. Additionally, if each regulator can reduce $\theta$ negative interactions, then the total number of negative interactions becomes $\rho\left(f-\eta c\right)\left(f-\eta c-1\right)-\theta r$ leading to the final utility function,
@@
where $\gamma_{2}$ is the unit cost of a regulator and $\gamma_{3}$ is the unit cost of a compartment (including the cost of negative interactions and regulators within a compartment). Here we assume that there is a proportional cost of $r$ and $c$ and a linear increase in benefit associated with $f$ such that $R=\gamma_{2}r$, $C=\gamma_{3}c$, and $B=bf$, where $b$ is the productive output of an average individual. All terms are measured in dollars or energy, depending on the system of interest. It is also useful to note that the total size of the system, $N$, which is often what is measured, is the sum of functional and regulatory components, $N=f+r$.
@@
Later we will show that compartments are not beneficial below a certain size, so we begin by considering an optimization for a system that only has regulators. As shown in the Supplementary Information, optimizing the utility function $\mathcal{L}$ yields the optimal number of regulators in this scenario:
$r_{\rm opt}\sim\frac{f^{2}\rho}{\theta}\;,$ (3)
@@
illustrating that $r$ grows like $f^{2}$ adjusted by the ratio of the probability of negative interactions, $\rho$, to the number of interactions that each regulator can mitigate, $\theta$. Equation 3 allows us to express the total size of the system in terms of $r$ alone: $N=f+r=\sqrt{\theta/\rho}\;r^{1/2}+r$. Note that this leads to two distinct regimes characterized by how $r$ relates to $N$: for small $f$ (which corresponds to small $r$, given Eq. 3), $r\propto N^{2}$, while for large $f$ (and correspondingly for large $r$), $r\propto N$.
The first case describes an ideal scenario, most closely approximated by bacteria. There are two ways in which segregation happens in cells, either through the physical separation of expressed proteins, or through the genomic groupings of genes that are co-expressed and/or co-regulated. Bacteria have genes that generally interact in an all-to-all environment of the cytoplasm, and they do not have genes that are segregated in different chromosomes. Thus, we expect bacteria to be closest to the quadratic scaling of systems without compartments. The literature report superlinear exponents as high as $1.86$ (see Supplementary Information for a review). However, bacteria are known to shift scaling toward the largest and smallest cells for a variety of biophysical and physiological reasons that lead to asymptotic behavior [kempes2012growth, kempes2016evolutionary, gondhalekar2023scaling, supo2025percolation]. It is also the case that large bacteria acquire effective compartments that occur in several ways. First, proteins are often transported to a specific location, such as the membrane to perform functions and this effectively isolates them from other proteins. Second, recently it has been shown that liquid-liquid phase separation produces a variety of membrane-less organelles in bacteria [jin2021membraneless]. Third, despite having a single chromosome, bacterial genomes are partially compartmentalized using operons (clusters of co-expressed genes), and higher-order modulons (groups of operons) [osbourn2009operons]. Thus, we should expect larger bacteria to deviate from quadratic scaling due to the addition of compartments. We perform a breakpoint analysis on the binned data [gondhalekar2023scaling] and find two distinct scaling relationships, where the first gives the exponent of $1.65$for small cells, and the second gives $1.37$ for the largest bacteria. It is important to note that the first scaling exponent excludes the smallest cells where only a small amount of data is available and where they demonstrate their own scaling regime (see Supplementary Information). The deviation of $1.65$ from quadratic is likely due to the genomic co-location of sets of functions. The deviation of $1.37$ from linear is likely due to the more imperfect and more informal compartments of large bacteria.
@@
For the second case, where $N\propto r$, there is a critical size at which regulators overwhelm the system causing $\mathcal{L}=0$, which occurs at $f_{r}=1+b\theta/(\gamma_{2}\rho)$. Here, $f_{r}$ is the maximum number of functional elements for a system with regulators but no compartments. This upper bound depends on the ratio of the benefit per functional element to an effective cost, namely, the unit cost of a regulator times the likelihood of a negative interaction. The ratio of $b\theta$ to $\gamma_{2}\rho$ could be quite large given that the probability of negative interactions could be small and the cost of an average individual, including regulators, $\gamma_{2}$, should be small relative to the productive output, $b$, of an average individual. In addition, a regulator may be able to handle many interactions so that $\theta$ could be much larger than 1. For example, if $\rho=0.10$, $\gamma_{2}=b/2$, and $\theta=10$, then $f\approx 200$, implying that an organization that handles negative interactions with only regulators could reach a reasonable size without introducing divisions or compartments.
@@
As discussed above, $f_{r}$ is the maximum number of functional elements that can operate optimally in a system with only regulators; beyond this number, compartments are needed. In Supplementary Information, we optimize the utility function $\mathcal{L}$ with respect to both number of compartments $c$, and number of regulators $r$, and find that the optimal number of compartments, $c_{\rm opt}$, is
$c_{\rm opt}\sim\frac{f}{\eta}$ (4)
@@
This scaling comes with a variety of cost and size requirements that we discuss in detail in the Supplementary Information. These results show that in addition to $c\sim f$ the optimal regulators follow $r\sim f$. Because of this linear scaling, the total number $N=f+r$ is also linear: $r\propto N$ and $c\propto N$. In this regime, the quadratic requirements of negative pairwise interactions are handled by a combination of compartments, and regulators in those compartments, each of which scales linearly. However, this solution only occurs after the transition where compartments become inexpensive enough to be viable.